# Neural Ordinary Differential Equations for Learning and Extrapolating System Dynamics Across Bifurcations

### v1

@@ 5 Results
For the primary experiment, to evaluate the Neural ODE’s ability to learn vector fields from trajectory data, we trained the model on the primary dataset consisting of ten time series. As shown in Fig. 5, the model is able to reconstruct vector fields from limited trajectory data, qualitatively capturing the underlying system dynamics. Notably, the model not only replicates the behaviour observed in the training trajectories but also accurately extrapolates the onset of a global bifurcation leading to system collapse. Since global bifurcations are characterized by large-scale changes in the state space that extend beyond local trajectories, this result indicates that the model has captured structural information beyond the training data.
A closer examination of Fig. 5 reveals that the Neural ODE successfully extracted several key dynamic components. First, the learned vector field exhibits two saddle nodes along the $x$-axis, positioned similarly as those in the true system. Second, along both the $x$- and $y$-axes, the component of the flow perpendicular to the respective axis is close to zero. The derivatives not reaching exactly zero can be attributed to the soft formulation of the physics-constraint. Overall, despite minor deviations, the model accurately reconstructs the qualitative behaviour of the system across different values of $\alpha$.
Figure 6: Reconstructed bifurcation diagrams for the different experiments. Each subfigure shows the true dynamics (black), best-performing learned dynamics (blue) and the learned dynamics for other random seeds (grey).
@@
We further evaluated the performance of the Neural ODE in different experiments and compared the results to those obtained in the primary experiment. Fig. 6 shows the bifurcation diagrams generated by the trained models for each experiment. Details on the simulation and visualisation process can be found in the Appendix C. The figure illustrates the model’s convergence behaviour. As in Fig. 2, limit cycles are presented by plotting the minimum and maximum value. To account for stochastic variability in initialisation and training, the model was trained using five different random seeds. The best-performing seed, based on MAE relative to the true system, is shown in blue.
When trained on time series that span both the stable focus and limit cycle regimes (regimes A and B), the model consistently predicts the occurrence of a global bifurcation across all random seeds. While the collapse is correctly captured, the system does not reach exactly zero due to the soft nature of the physics-constraint. Notably, the zero steady state is also absent from the learned dynamics in the regime on which the model was trained. This can be explained by the lack of time series in the dataset that converge to zero, making it inherently difficult for the model to learn and reproduce a state that is not represented in the training data.
When trained exclusively on time series from the limit cycle regime (regime B), the model shows varied performance across different random seeds. The best-performing seed is able to predict both the global bifurcation on the right and the transition to a stable focus on the left. However, results from other seeds are less consistent, highlighting the limited coverage of the training data and the model’s resulting uncertainty when extrapolating beyond the observed regime.
When trained solely on data from the stable focus regime (regime A), the model consistently predicts a transition to limit cycle behaviour across all random seeds, but fails entirely to capture the global bifurcation. This suggests that the training data lacks sufficient information to fully reconstruct key structures of the state space, such as saddle points along the $x$-axis, which play a crucial role in shaping the bifurcation dynamics.
@@
To assess how the amount of training data influences the model performance, for training we used datasets with limited data compared to the primary experiment. In one case we decreased the number of sampled $\alpha$-values. This reduction introduced a varied performance for the different random seeds, notably also within the training regime. While the best-performing seed still produced a qualitatively accurate reconstruction of the bifurcation structure, other seeds showed less consistent results. Importantly, however, all random seeds correctly predicted a global collapse, indicating that the model retains some robustness in identifying critical system transitions even with limited parameter diversity.
In the other case, we restricted the number of initial conditions per $\alpha$-value. Despite this limitation, the model generally succeeded in predicting limit cycle dynamics and the transition leading to collapse. However, the predicted global bifurcation occurred at a lower $\alpha$-value than in the primary experiment.
For both types of data reduction, the best-performing random seeds were able to reproduce the qualitative behaviour of the true system. Nonetheless, a visual comparison with results from the primary experiment reveals a noticeable decline in quantitative accuracy (Appendix C), suggesting that while the overall dynamics remain consistent, the precise locations and timing of bifurcations become less reliable with less training data.
@@
Finally, we introduced different levels of measurement noise into the training data. At low noise levels, the predicted bifurcation parameter range and amplitude of limit cycles showed considerable variation across different random seeds. Nevertheless, the best-performing seed still produced a reasonable qualitative approximation of the system dynamics. With higher noise levels, the model’s predictions for the onset of limit cycles became less precise, exhibiting larger deviations compared to the primary, noise-free experiment. Despite this increased uncertainty, the model remained capable of distinguishing between distinct dynamical regimes, successfully identifying both the emergence of limit cycles and the eventual occurrence of the global bifurcation.
@@ 6 Discussion
In this study, we demonstrated that a Neural ODE model can successfully uncover the underlying bifurcation structure of a predator-prey system training only on limited trajectory data. Results from the experiments (Fig. 4 and 5) indicate that the model is capable of forecasting global bifurcations beyond the parameter regions represented in the training data. In Experiment 1, we restricted training data to specific regimes of the bifurcation diagram. The model’s ability to extrapolate varied depending on the chosen regime. When training trajectories did not sufficiently span the state space in order for the model to capture key dynamical features, such as saddle points, the Neural ODE had limited success in predicting transitions to other regimes. In Experiment 2, we reduced the quantity of training data compared to the primary experiment either by sampling fewer parameter values or by limiting the number of initial conditions. Although the best-performing random seeds still captured the system’s qualitative behaviour, performance became less consistent across seeds, suggesting increased model uncertainty. Finally, in Experiment 3, we introduced measurement noise into the primary dataset. Despite this added disturbance, the model retained its ability to extrapolate distinct dynamical regimes, demonstrating a certain level of robustness to noisy data.
@@
While our results highlight the potential of Neural ODEs to learn entire vector fields directly from trajectory data, they also reveal a key challenge: the quality of extrapolation depends on the information that can be inferred from the training data. When the training data spans a limited part of the state space and essential aspects, such as saddle points, can not be inferred, the model lacks the basic components to reconstruct the full bifurcation structure. This indicates that extrapolation improves when the training data offers more extensive coverage of the system’s possible dynamic behaviours.
@@
In real-world systems, however, it is often difficult or impossible to observe the full range of system behaviour, since features like unstable fixed points, are often not represented by the data. In such cases, integrating prior knowledge or mechanistic constraints into the learning process, e.g. through physics-informed neural networks or data augmentation using synthetic data, could enhance a model’s ability to generalise to unobserved regions[42, 43, 44, 45]. We noted a similar effect in our own work: incorporating a physics-based loss led to improved model performance, since it gave the model valuable information about the state space.
@@
In our paper, we kept the bifurcation parameter $\alpha$ constant within each trajectory to enable more interpretable analysis of the model’s extrapolation behaviour. Allowing $\alpha$ to vary dynamically would effectively place trajectories in a higher-dimensional space, making the dynamics more complex to visualise and interpret. Starting with constant parameters therefore provided a natural and informative first step before extending the approach to scenarios involving time-dependent bifurcation parameters. While our experimental setup was deliberately simplified for clarity, the method itself is not limited to this scenario. It is broadly applicable to any continuous dynamical system for which trajectory data is available. In particular, the Neural ODE framework can incorporate dynamically-changing parameters with minimal modification, making it well-suited for real-world systems where $\alpha$ is generally known but varies over time.
@@
In this study, we used a basic Neural ODE architecture to model the system dynamics. While this basic formulation already demonstrated strong potential, a range of extensions to the Neural ODE framework have been proposed [46, 47, 27, 32, 48, 49], including Augmented Neural ODEs [50], which may offer improved accuracy in capturing complex dynamical behaviours or greater robustness in the presence of noise. A promising direction for future work would be to systematically compare the performance of standard Neural ODEs, their extensions, and discrete-time machine learning methods in tasks such as forecasting global bifurcations and reconstructing bifurcation diagrams from data. While such comparisons are not straight-forward, especially since discrete-time models typically lack access to the underlying vector field and operate at fixed time steps, they could provide valuable insight into the trade-offs between continuous and discrete approaches in terms of accuracy, interpretability, and computational cost.
@@
While the above comparisons focus on different machine learning models for forecasting and reconstruction, another valuable perspective is to examine how the proposed approach compares to, and could potentially be integrated with, alternative frameworks for predicting critical transitions. A common approach to predicting critical transitions, specifically local bifurcations, is through early warning signals based on slowed recovery after small perturbations [51, 52, 53]. Unfortunately, these signal are not observed before all tipping points, including global bifurcations [54, 55, 56]. Furthermore, because these signals do not model or capture the system’s dynamics, they cannot be used to extrapolate behaviour beyond the bifurcation point [57]. This limitation points to a potential integration with Neural ODEs. Since early warning signals typically rely on process noise to reveal subtle changes in stability, introducing similar noise-driven variability into the training data could help the Neural ODE detect potential stability changes in trajectory data more reliably. In future work, these signals could potentially be treated as an additional input to improve the model’s ability to anticipate and reconstruct complex bifurcations.
@@
In conclusion, this work demonstrates that Neural ODEs can be used not only to model continuous-time dynamics from trajectory data, but also to recover underlying bifurcation structures, even in the presence of global bifurcations. By explicitly including the bifurcation parameter and learning the full vector field, our approach offers a flexible and interpretable framework for studying critical transitions in dynamical systems. These results motivate future extensions to more realistic systems, including those with multiple or time-varying drivers and unknown environmental interactions.
@@ Acknowledgements
Appendix A Data overview
In this section, we provide additional details about the data used to train the models for the various experiments. Table 1 summarizes the initial conditions and bifurcation parameter values that were used to generate the time series data. For each configuration, we used numerical integration to simulate the solution trajectories and obtained time series data by sampling at discrete time steps $t={0,1,2,\dots,100}$. Note that in each of the experiments one of the datasets overlaps with the primary experiment. This was done for easy comparison.
Fig. 7 provides an example of the trajectory data for a specific value of $\alpha$. To highlight the impact of the added measurement noise, the figure also displays the corresponding time series under different noise levels, illustrating how the original dynamics become increasingly distorted as the noise level increases.
Table 1: Overview of experimental setup with $\alpha$-values, initial conditions, and noise levels for the different datasets. Differences with respect to the primary experiment are indicated in cursive.
@@ Appendix A Data overview
[…]
@@
[b]
@@
Appendix B Hyperparameters and tuning
We tuned our model hyperparameters once for the primary experiment. To perform hyperparameter tuning, we split our primary dataset into a training and validation set, ensuring that the validation set contained a value of $\alpha$ outside the range seen during training to specifically test the model’s ability to extrapolate. More specifically, from the primary dataset we selected the two trajectories with $\alpha=0.62$ for validation, while the remaining eight trajectories were used for training.
Hyperparameter selection was conducted via a full grid search over the ranges specified in Table 2. Our aim was not to optimize model performance, but simply to demonstrate that Neural ODEs have the potential to learn bifurcation structures, so we explored only a limited range of tuning parameters. Each configuration was evaluated across three different random seeds to account for stochastic variability in model initialization and training. For each trial, we computed the following MAE on the validation set:
[…]
where $\mathbf{z}_{i}=[x_{i},y_{i}]$ are our validation data points, $\tilde{\mathbf{z}}_{i}$ our predicted state variables, and $N_{val}$ is the number of validation data points. Note that the error on the validation set is calculated differently from the loss function in Eq. 4, since the physics-loss is not taken into account. The hyperparameter combination with the lowest average MAE across the different seeds was selected and used to train the models for all experiments.
Table 2: Hyperparameters considered during tuning, along with their tested values. Best performing hyperparameter combination is indicated in bold (lowest average MAE over three different random seeds).

### v2

@@ 4 Results
Figure 4: Lorenz system trajectories. Top: True system. Bottom: Dynamics learned by the Neural ODE.
Figure 5: Bifurcation diagram for Lorenz system. Top: True system. Bottom: Dynamics learned by the Neural ODE.
From the Neural ODE, we are able to obtain bifurcation structures by simulating the dynamics of the systems. The results are presented in this section. Details on the simulation and visualization process can be found in Appendix D.
Test case I: Lorenz system
For the Lorenz system, we trained the Neural ODE on non-chaotic behaviour and evaluated its ability to extrapolate into the chaotic regime. Fig. 4 shows three-dimensional solution trajectories for multiple values of $r$. The learned trajectories closely match the true trajectories, particularly within the training range. More interestingly, the Neural ODE is able to extrapolate chaotic behaviour beyond the training data. While the extrapolated chaotic trajectories do not exactly replicate the true dynamics, they accurately capture the onset and qualitative features of chaos.
A similar pattern is observed in the bifurcation diagram shown in Fig. 5. The Neural ODE successfully learns the two stable fixed-point branches present in the training regime and, more importantly, extrapolates the emergence of chaotic behaviour beyond the training range. The value of $r$ at which chaos first seems to appears in the learned system closely matches that of the true Lorenz system, indicating that the model captures not only the qualitative dynamics but also reasonably approximates the bifurcation point.
To the left of the training regime, however, the extrapolation is less accurate: the two stable fixed-point branches are slightly distorted compared to the true dynamics, reflecting minor limitations in the model’s predictive accuracy farther outside the training range.
Test case II: Rössler system
For the Rössler system, we trained the Neural ODE on chaotic behaviour and evaluated whether it could learn the underlying dynamics sufficiently to reproduce limit cycles and period-doubling bifurcations occurring at lower values of the bifurcation parameter $c$. Remarkably, despite being trained only on chaotic trajectories, the Neural ODE is able to recover non-chaotic behaviour in the extrapolated regime. The bifurcation diagram in Fig. 7 shows that the model captures not only the chaotic behaviour present in the training data, but also reproduces the limit cycles and period-doubling structure to the left of this regime with reasonable accuracy.
Closer to the training regime, the learned dynamics more closely resemble the true system. However, the exact parameter values at which period-doubling bifurcations occur are slightly shifted compared to the true system. Interestingly, the Neural ODE also predicts an additional period-doubling bifurcation around $c\approx 3.35$, which is not present in the true system. These results highlight both the model’s impressive ability to generalize beyond the training data, but also the subtle limitations of its extrapolation.
Figure 7: Bifurcation diagram for Rössler system. Top: True system. Bottom: Dynamics learned by the Neural ODE.
Test case III: Predator-prey system
Figure 8: Reconstructed bifurcation diagrams for the predator-prey system for the different experiments. Each subfigure shows the true dynamics (black), best-performing learned dynamics (blue) and the learned dynamics for other random seeds (grey).
@@
For the predator-prey system, the model is able to reconstruct vector fields from limited trajectory data, qualitatively capturing the underlying system dynamics, as shown in Fig. 6. Notably, the model not only replicates the behaviour observed in the training trajectories but also extrapolates the onset of a global bifurcation leading to system collapse. Since global bifurcations are characterized by large-scale changes in the state-space that extend beyond local trajectories, this result indicates that the model has captured structural information beyond the training data. A closer examination reveals that the Neural ODE successfully extracts several key dynamic components. In particular, along both the $x$- and $y$-axes, the component of the flow perpendicular to the respective axis remains close to zero. Overall, despite minor deviations, the model accurately reconstructs the qualitative behaviour of the system across different values of $\alpha$.
@@
We further evaluated the performance of the Neural ODE in different experiments and compared the results to those obtained in the primary experiment. Fig. 8 shows the bifurcation diagrams generated by the trained models for each experiment. To account for and showcase the stochastic variability in initialisation and training, the model was trained using five different random seeds. Each of these is displayed in the figure with the best-performing seed, based on MAE relative to the true system, shown in blue. (See Appendix D for calculation.)
When trained on both the stable coexistence and limit cycle regimes A and B, the model consistently predicts the global bifurcation across all seeds, though the collapse does not reach exactly zero due to the soft nature of the physics-constraint. Training only on the limit cycle regime B yields variable performance: the best seeds capture both the global bifurcation and the transition to a stable focus, while others fail, highlighting the limited coverage of the training data and the model’s resulting uncertainty when extrapolating beyond the observed regime. When trained solely on the stable focus regime A, the model consistently predicts a transition to limit cycles but fails entirely to capture the global bifurcation. This suggests that the training data lacks sufficient information to fully reconstruct key structures of the state-space, such as saddle nodes along the $x$-axis, which play a crucial role in shaping the bifurcation dynamics.
Notably, the zero steady state is absent from the learned dynamics, even in the parameter range it was trained on. This can be explained by the lack of time series in the dataset that converge to zero, making it inherently difficult for the model to learn and reproduce a state that is not represented in the training data.
@@
To assess how the amount of training data influences the model performance, we trained on datasets with reduced coverage compared to the primary experiment, either by decreasing the number of sampled $\alpha$-values or by limiting the number of initial conditions per $\alpha$-value. Fewer $\alpha$-values led to varied performance across random seeds, even within the training regime; while the best seeds qualitatively reconstructed the bifurcation structure, others were less consistent. Nevertheless, all seeds correctly predicted a global collapse, indicating robustness in identifying critical transitions despite reduced parameter diversity. Limiting initial conditions generally preserved the model’s ability to predict limit cycle dynamics and collapse, though the global bifurcation appeared at a lower $\alpha$-value than in the primary experiment.
For both types of data reduction, the best-performing random seeds were able to reproduce the qualitative behaviour of the true system. Nonetheless, a visual comparison with results from the primary experiment reveals a noticeable decline in quantitative accuracy (Appendix E), suggesting that while the overall dynamics remain consistent, the precise locations and occurrences of bifurcations become less reliable with less training data.
@@
Finally, we introduced different levels of measurement noise into the training data. At low noise levels, the predicted bifurcation locations and amplitude of limit cycles showed considerable variation across different random seeds. Nevertheless, the best-performing seed still produced a reasonable qualitative approximation of the system dynamics. With higher noise levels, the model’s predictions for the onset of limit cycles became less precise, exhibiting larger deviations compared to the primary, noise-free experiment. Despite this increased uncertainty, the model remained capable of distinguishing between distinct dynamical regimes, successfully identifying both the emergence of limit cycles and the eventual occurrence of the global bifurcation. Further discussion of the quantitative accuracy can be found in Appendix E.
@@ 5 Discussion
In this study, we demonstrated that a Neural ODE model can successfully uncover underlying bifurcation structures, training only on limited trajectory data. For the Lorenz system, the model extrapolated chaotic dynamics from non-chaotic training data. For the Rössler system, training on chaotic data enabled qualitatively accurate prediction of non-chaotic behaviour, including period-doubling bifurcations. Experiments on the predator-prey system indicated that Neural ODEs are capable of forecasting global bifurcations beyond the parameter regions represented in the training data. Reducing the amount or quality of training data for this test case decreased consistency across random seeds, while added noise introduced variability but did not remove the model’s ability to recover qualitative regime changes.
@@
While our results highlight the potential of Neural ODEs to learn vector fields directly from trajectory data, they also reveal a key challenge: the quality of extrapolation depends on the information that can be inferred from the training data. When the training data spans a limited part of the state-space and essential aspects, such as saddle nodes, can not be inferred, the model lacks the basic components to reconstruct the full bifurcation structure. This indicates that extrapolation improves when the training data offers more extensive coverage of the system’s possible dynamic behaviours.
@@
In real-world systems, however, it is often difficult or impossible to observe the full range of system behaviour, since features like unstable fixed points, are often not represented by the data. In such cases, integrating prior knowledge or mechanistic constraints into the learning process, e.g. through physics-informed neural networks or data augmentation using synthetic data, could enhance a model’s ability to generalise to unobserved regions[45, 46, 47, 48]. We noted a similar effect in our own work: incorporating a physics-based loss for the predator-prey system led to improved model performance, since it gave the model valuable information about the state-space.
@@
In our paper, we kept the bifurcation parameters constant within each trajectory to enable more interpretable analysis of the model’s extrapolation behaviour. Allowing bifurcation parameters to vary dynamically would effectively place trajectories in a higher-dimensional state-space, making the dynamics more complex to visualise and interpret. Starting with constant parameters therefore provided a natural and informative first step before extending the approach to scenarios involving time-dependent bifurcation parameters. While our experimental setup was deliberately simplified for clarity, the method itself is not limited to this scenario. It is broadly applicable to any continuous dynamical system for which trajectory data is available. In particular, the Neural ODE framework can incorporate dynamically-changing parameters with minimal modification, making it well-suited for real-world systems with external drivers that vary over time.
@@
In this study, we used a basic Neural ODE architecture to model the system dynamics. While this basic formulation already demonstrated strong potential, a range of extensions to the Neural ODE framework have been proposed [49, 50, 27, 51, 52, 53], including Augmented Neural ODEs [54], which may offer improved accuracy in capturing complex dynamical behaviours or greater robustness in the presence of noise. Compared to discrete-time approaches such as Recurrent Neural Networks (RNNs) or Long Short-term Memory Networks (LSTMs), Neural Ordinary Differential Equations naturally operate in continuous time and directly approximate the underlying vector field, which can be advantageous when analyzing bifurcations. A promising direction for future work would be to systematically compare the performance of standard Neural ODEs, their extensions, and discrete-time machine learning methods in tasks such as forecasting global bifurcations and reconstructing bifurcation diagrams from data. While such comparisons are not straight-forward, especially since discrete-time models typically lack access to the underlying vector field and operate at fixed time steps, they could provide valuable insight into the trade-offs between the continuous and discrete frameworks in terms of accuracy, interpretability, and computational cost.
@@
While the above comparison focusses on different machine learning models for forecasting and reconstructing bifurcations, another valuable perspective is to examine how the proposed approach compares to, and could potentially be integrated with, alternative frameworks for predicting critical transitions. A common approach to predicting critical transitions, specifically local bifurcations, is through early warning signals based on slowed recovery after small perturbations [55, 56, 57]. Unfortunately, these signals are not observed before all tipping points, including global bifurcations [58, 59, 60]. Furthermore, because these signals do not model or capture the system’s dynamics, they cannot be used to extrapolate behaviour beyond the bifurcation point [61]. This limitation points to a potential integration with Neural ODEs. Since early warning signals typically rely on process noise to reveal subtle changes in stability, introducing similar noise-driven variability into the training data could help the Neural ODE detect potential stability changes in trajectory data more reliably. In future work, these signals could potentially be treated as an additional input to improve the model’s ability to anticipate and reconstruct complex bifurcations.
@@
In conclusion, this work demonstrates that Neural ODEs may be used not only to capture system dynamics from trajectory data, but also to recover underlying bifurcation structures, such as heteroclinic bifurcations, chaos and period-doubling. By explicitly including the bifurcation parameter and learning the full vector field, our approach offers a flexible and interpretable framework for studying critical transitions in dynamical systems. These results motivate future applications to more realistic systems, including those with multiple or time-varying parameters and unknown environmental interactions.
@@ Acknowledgements
Appendix A System analysis predator-prey
In this appendix we provide more details on the system dynamics of the predator-prey system. Fig. 9 illustrates the bifurcation structure of the system, highlighting the different dynamical regimes. Note that we limit ourself to the bifurcation parameter range $\alpha\in(0.6,0.8)$, since the zero steady state is only stable for $\alpha>0.6$. Within this range the system exhibits three distinct dynamical regimes, which allows us to analyse if the model is able to extrapolate transitions between qualitatively different behaviours. The system has four steady states: the zero steady state, two saddle nodes on the $x$-axis and a coexistence steady state.
[…]
In Regime A the coexistence steady state is a stable focus, meaning that for a range of initial conditions the predator and prey populations settle into a steady coexistence. At $\alpha=0.67$, the system undergoes a Hopf bifurcation, causing the coexistence steady state to lose stability, giving rise to stable limit cycles, which characterise Regime B. These limit cycles persist until $\alpha\approx 0.71$, where the system undergoes a global heteroclinic bifurcation; the limit cycle collides with the two saddle nodes, leading to a collapse. In Regime C ($\alpha>0.71)$, the only remaining stable equilibrium is the zero steady state, meaning that both populations eventually go extinct, marking a critical transition point, that has previously been interpreted as an ecological phenomenon referred to as over-exploitation[62].
Appendix B Data generation
In this section, we provide additional details about the data used to train the models for the various systems. Table 1 summarizes the initial conditions and bifurcation parameter values that were used to generate the time series data. For the Lorenz and Rössler systems, we normalised the data using standard normalisation, whereas for the predator–prey model, the data was left unnormalised.
Table 1: Overview of experimental setup for different systems. Bifurcation parameter values, initial conditions, and noise levels for the different datasets. The values of the bifurcation parameter that were used as the validation set are indicated in grey. Predator-prey system: differences with respect to the primary experiment are indicated in cursive.
@@ Appendix B Data generation
Test case I: Lorenz $r$ $(x,y,z)$
5.0, 7.5, 10, 12.5, 15, (1,1,1),(-1,-1,1)
17.5, 20, 22.5
Test case II: Rössler $c$ $(x,y,z)$
4.5, 4.6, 4.7, 4.8, 4.9 (-4, -4, 0)
5.0, 5.1, 5.2, 5.3, 5.4
Test case III: Predator-prey $\alpha$ $(x,y)$ $\sigma$
@@
Test case I: Lorenz
We used numerical integration of the Lorenz system to simulate the trajectories for the given values of the bifurcation parameters and initial conditions. Starting from time $t=0$, we sampled 100 time points with time step $dt=0.05$.
Test case II: Rössler
For the Rössler system, we similarly used numerical integration to simulate the trajectories. We wanted the system already to be converged to an attractor, so we used a warm-up time of $t=475$ and them sampled 200 time points with a time step $dt=0.125$.
Test case III: Predator-prey
For the predator-prey system, for each configuration illustrated in Table 1, we used numerical integration to simulate the solution trajectories and obtained time series data by sampling at discrete time steps $t={0,1,2,\dots,100}$. Note that in each of the experiments one of the datasets overlaps with the primary experiment. This was done for easy comparison.
Fig. 10 provides an example of the trajectory data for a specific value of $\alpha$. To highlight the impact of the added measurement noise, the figure also displays the corresponding time series under different noise levels, illustrating how the original dynamics become increasingly distorted as the noise level increases
@@
Appendix C Training and tuning procedure
C.1 Loss function
The loss function for the Lorenz and Rössler test cases was defined using the MAE:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}{\|\mathbf{z}_{i}-\tilde{\mathbf{z}}_{i}\|}_{1},$ (7)
where $\mathbf{z}_{i}=[x_{i},y_{i}]$ are our training data points, $\tilde{\mathbf{z}}_{i}=[\tilde{x}_{i},\tilde{y}_{i}]$ are our predicted state variables and $N$ is the number of data points. For the predator-prey system, to guide the training process, we combined a MAE with a physics-informed term to obtain our loss function:
$\mathcal{L}=$ $\frac{1}{N}\sum_{i=1}^{N}{\|\mathbf{z}_{i}-\tilde{\mathbf{z}}_{i}\|}_{1}$
$+\lambda\left[\frac{1}{M}\sum_{j=1}^{M}\left|\frac{d\tilde{x}}{dt}(0,y_{j})\right|+\right.\left.\frac{1}{M}\sum_{j=1}^{M}\left|\frac{d\tilde{y}}{dt}(x_{j},0)\right|\right],$
where $M$ indicates the number of points on which we evaluate the physics-loss. The physics part of the loss function enforces that the axes are invariant by penalizing non-zero derivatives of the system along specific axes. In our experiments, adding the physics-informed term improved generalisation to unseen parameter ranges by guiding the model training in regions of the vector field not covered by the data. The hyperparameter $\lambda$ is the physics-loss weight and determines how much the physics-loss contributes to the total loss. Note that the physics-informed term is not a hard constraint, but a soft constraint driving the derivatives on the axes towards zero.
C.2 Batching
C.3 Hyperparameter tuning
To perform hyperparameter tuning, we split our data for each of the systems into a training and validation set, ensuring that the validation set contained a value of the bifurcation parameter outside the range seen during training to specifically test the model’s ability to extrapolate. The trajectories that were used as validation are highlighted in Table 1 in grey. We conduct a random grid search (Lorenz and Rössler) or full grid search (predator-prey) and for each selected parameter combination we evaluated the model across three different random seeds to account for stochastic variability in model initialization and training. For each trial, we computed the following MAE on the validation set:
$\mathit{MAE}_{val}=\frac{1}{N_{val}}\sum_{i=1}^{N_{val}}{\|\mathbf{z}_{i}-\tilde{\mathbf{z}}_{i}\|}_{1},$ (9)
where $\mathbf{z}_{i}=[x_{i},y_{i}]$ are our validation data points, $\tilde{\mathbf{z}}_{i}$ our predicted state variables, and $N_{val}$ is the number of validation data points. The investigated and selected hyperparameters for each of the systems are listed in Table 2.
For the predator-prey system multiple experiments were done with different datasets, however we tuned our model parameters only once for the primary experiment. The hyperparameters remained fixed across all other experiments. In addition to the tuned hyperparameters, several other parameters were held fixed throughout all experiments. These are listed in Table 3.
Table 2: Hyperparameters considered during tuning, along with their tested values. Best performing hyperparameter combinations are indicated in bold (lowest average MAE over three different random seeds).